OpenStudy (anonymous):

Pwease help mehh Equivalent representation of 4^-4 ?

OpenStudy (aum):

\[ 4^{-4} = \frac{1}{4^4} = \frac{1}{256} \]

OpenStudy (anonymous):

Oh god, I forgot the fraction thing to make the exponent a positive number! Thank you so much!!!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

Do you mind helping me with another?

OpenStudy (anonymous):

\[(6^4)^6\div6^24\] The 4 on the very end of this equation is part of the exponent..

OpenStudy (aum):

\[\large \frac{(6^4)^6}{(6^2)^4} = \frac{6^{24}}{6^{8}} = 6^{24-8} = 6^{16} \]